# Cost-Effective, High-Performance Open-Source LLMs via Optimized Context Retrieval

### v2

@@ Abstract
Large Language Models (LLMs) in healthcare promise transformation, yet adoption is limited by concerns over factual accuracy and the high cost of proprietary models. This study demonstrates that optimized context retrieval unlocks cost-effective, high-performance healthcare AI using open-source LLMs, achieving a significantly improved cost-accuracy Pareto frontier for medical question answering and showcasing that open models can rival proprietary systems at a fraction of the cost. A key contribution is OpenMedQA, a novel benchmark for open-ended medical question answering that overcomes the limitations of multiple-choice formats—formats that we show lead to performance degradation in open-ended settings and often lack clinical realism. Further contributions include: (1) practical guidelines for implementing optimized context retrieval; (2) empirical validation of enhanced cost-effectiveness via the improved Pareto frontier; (3) the introduction of OpenMedQA for rigorous evaluation of open-ended medical QA; and (4) the release of prompt_engine alongside CoT/ToT/Thinking databases as community resources for cost-effective healthcare AI. Advancing optimized retrieval and open-ended QA benchmarking, we pave the way for more accessible and impactful LLM-powered healthcare solutions. All the materials have been made public here.
@@
Foundational models Prompt engineering Medical benchmarks Pareto Frontier Open Source Cost-Effectiveness
@@ 1 Introduction
Our work offers a practical pathway toward democratizing access to high-quality healthcare AI, shifting the focus from solely relying on ever-larger, more expensive models to leveraging efficient, cost-optimized architectures and techniques. Contributions of this work are:
@@ 2 Related Work
Addressing the critical challenge of factuality in LLMs has become a central topic in AI research. This is particularly the case within high-stakes domains, like healthcare, where reliability is paramount and misinformation can have severe consequences. Initial efforts to enhance LLM accuracy have centered on In-Context Learning (ICL) techniques [1]. These methods, applied at inference time without reliance on external data sources, encompass sophisticated prompting strategies such as Chain of Thought (CoT) [25], Tree of Thought (ToT) [26], Self-Consistency (SC) [24]and Reasoning models [3]. While these techniques demonstrably improve response coherence and performance on intricate tasks, they often fall short of guaranteeing the stringent degree of factuality demanded in critical medical applications.
@@
In the last few years, a performance gap between large proprietary LLMs and their open-weight counterparts [12, 6] has persisted. However, recent progress in open-source LLM development [13, 4, 3], combined with the potential of efficient RAG systems, suggests a promising pathway toward closing this gap while addressing the critical imperative of cost-effectiveness. In this regard, Pareto frontiers provide a valuable framework for analyzing cost-accuracy trade-offs, [17], enabling a more nuanced assessment by explicitly considering computational efficiency as a key performance indicator.
To comprehensively evaluate LLM performance in healthcare, it is crucial to acknowledge the inherent limitations of popular Multiple-Choice Question Answering (MCQA) benchmarks. MCQA benchmarks present significant shortcomings in their ability to reflect the multifaceted complexities of real-world clinical scenarios [6]. Authentic clinical practice predominantly involves answering open-ended questions, demanding nuanced, comprehensive, and contextually rich responses that extend far beyond the constraints of pre-defined options. While CoT prompting has been explored for open-ended QA tasks, research in this critical area remains comparatively limited and non-exhaustive [12, 20].
@@ 4.1 SC-CoT Experiments
SC-CoT’s performance was then analyzed in relation to ensemble size ($N$), balancing accuracy gains with computational cost. Here, $N$ represents the number of responses generated per question, with the final answer determined by majority voting among the selected options. Such trade-off can be seen in Figure 3, where performance as well as the footprint consistently have an upward trend with the first five ensembles yielding around 3.5% accuracy gains. However, the additional ensembles provide minimal performance improvements at higher computational costs (up to 1.76kg of $CO_{2}$ for 25 ensembles). Based on this, 5 ensembles were used for all the experiments unless specified otherwise.
Figure 3: Accuracy vs. $CO_{2}$ emissions for increasing ensemble sizes in the SC-CoT setting. The solid lines represent accuracy trends for each dataset, while the dashed black line indicates the average accuracy. The shaded bars show $CO_{2}$ emissions in kg, highlighting the trade-off between performance gains and environmental cost.
@@ 4.3 State-of-the-art Comparison
Most importantly, the results demonstrate that optimized context retrieval provides a powerful pathway to achieve cost-effective high performance in healthcare AI. This cost reduction is visually and quantitatively substantiated by the Pareto frontier analysis presented in Figure 1, which demonstrably illustrates the shift toward a more efficient accuracy-cost trade-off enabled by the approach. By leveraging open-source models like Aloe-Beta-70B and DeepSeek-R1 in conjunction with optimized CR, state-of-the-art performance is achieved while drastically reducing operational costs. The cost was estimated by multiplying the total tokens with costs from the most affordable provider as of date  [^1] Artifiical Analysis: https://artificialanalysis.ai/, while the values for the closed models were taken from [17]. This finding has significant practical implications for democratizing access to high-quality healthcare AI, particularly in resource-constrained environments, where cost-effectiveness is a paramount consideration.

### v3

# Pareto-Optimized Open-Source LLMs for Healthcare via Context Retrieval

@@ Abstract
This study leverages optimized context retrieval to enhance open-source Large Language Models (LLMs) for cost-effective, high performance healthcare AI. We demonstrate that this approach achieves state-of-the-art accuracy on medical question answering at a fraction of the cost of proprietary models, significantly improving the cost-accuracy Pareto frontier on the MedQA benchmark. Key contributions include: (1) OpenMedQA, a novel benchmark revealing a performance gap in open-ended medical QA compared to multiple-choice formats; (2) a practical, reproducible pipeline for context retrieval optimization; and (3) open-source resources (Prompt Engine, CoT/ToT/Thinking databases) to empower healthcare AI development. By advancing retrieval techniques and QA evaluation, we enable more affordable and reliable LLM solutions for healthcare. All the materials have been made public here.
@@
Open-Source LLMs Healthcare AI Context Retrieval OpenMedQA Pareto Frontier Cost-Efficiency Medical Question Answering
@@ 1 Introduction
Our work offers a practical pathway toward reducing barriers to high-quality healthcare AI, shifting the focus from solely relying on ever-larger, more expensive models to leveraging efficient, cost-optimized architectures and techniques. Contributions of this work are:
@@ 2 Related Work
Ensuring factuality in LLMs is a key focus in AI research, especially in healthcare, where reliability is critical [6]. Initial efforts to enhance LLM accuracy have centered on In-Context Learning (ICL) techniques [1]. These methods, applied at inference time without reliance on external data sources, encompass sophisticated prompting strategies such as Chain of Thought (CoT) [25], Tree of Thought (ToT) [26], Self-Consistency (SC) [24]and Reasoning models [3]. While these techniques demonstrably improve response coherence and performance on intricate tasks, they often fall short of guaranteeing the stringent degree of factuality demanded in critical medical applications.
@@
In the last few years, a performance gap between large proprietary LLMs and their open-weight counterparts [12] has persisted. However, recent progress in open-source LLM development [13, 4, 3], combined with the potential of efficient RAG systems, suggests a promising pathway toward closing this gap while addressing the critical imperative of cost-effectiveness.
In this regard, Pareto frontiers provide a valuable framework for analyzing the inherent trade-off between model accuracy and computational cost [17]. Analyzing these frontiers requires quantifying the ’cost’ axis, which reflects the resources needed for deployment. This cost can encompass factors like API calls, inference time, or computational effort. Estimating computational effort, often proxied by metrics like floating-point operations (FLOPs) or energy consumption, allows for a more direct comparison of the resource efficiency of different models and configurations. Calculating this efficiency dimension is crucial for assessing the practical deployability and scalability of LLM solutions in real-world healthcare settings, potentially impacting equitable access across diverse institutions.
Finally, the predominant reliance on Multiple-Choice Question Answering (MCQA) benchmarks for evaluation presents limitations [6]. Real-world clinical practice necessitates handling open-ended questions requiring nuanced, generated responses. While CoT prompting has been explored for open-ended QA tasks, research in this critical area remains comparatively limited and non-exhaustive [12, 20].
@@ 4.1 SC-CoT Experiments
We then analyzed the impact of ensemble size ($N$) in SC-CoT. Here, $N$ represents the number of responses generated per question, with the final answer determined by majority voting among the selected options. As discussed in Section 2, evaluating cost-effectiveness requires analyzing the trade-off between accuracy gains and the associated computational resources. To quantify this for the ensemble size parameter, Figure 3 plots accuracy against Estimated Energy Consumption (kWh). This metric serves as a proxy for the computational effort and associated operational cost, estimated based on model size, token throughput, and typical GPU power draw profiles during inference. Analyzing this relationship allows us to identify an optimal balance for $N$.
This trade-off can be seen in Figure 3, where performance as well as the footprint consistently have an upward trend with the first five ensembles yielding around 3.5% accuracy gains. However, the additional ensembles provide minimal performance improvements at higher energy costs. Based on this, 5 ensembles were used for all the experiments, except in Table 4 where $N$ was set to 20 for maximal performance gains.
Figure 3: Accuracy vs. $CO_{2}$ emissions for increasing ensemble sizes in the SC-CoT setting. The solid lines represent accuracy trends for each dataset, while the dashed black line indicates the average accuracy. The shaded bars show power consumption in kWh, highlighting the trade-off between performance gains and environmental cost.
@@ 4.3 State-of-the-art Comparison
Most importantly, the results demonstrate that optimized context retrieval provides a powerful pathway to achieve cost-effective high performance in healthcare AI. This cost reduction is visually and quantitatively substantiated by the Pareto frontier analysis presented in Figure 1, which demonstrably illustrates the shift toward a more efficient accuracy-cost trade-off enabled by the approach. By leveraging open-source models like Aloe-Beta-70B and DeepSeek-R1 in conjunction with optimized CR, state-of-the-art performance is achieved while drastically reducing operational costs. The cost was estimated by multiplying the total tokens with costs from the most affordable provider as of date  [^1] Artifiical Analysis: https://artificialanalysis.ai/, while the values for the closed models were taken from [17]. This finding has significant practical implications for enhancing accessibility access to high-quality healthcare AI, particularly in resource-constrained environments, where cost-effectiveness is a paramount consideration.